# GVE-LPA: Fast Label Propagation Algorithm (LPA) for Community Detection in Shared Memory Setting

### 14 Dec 2023 v2

@@ Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for this purpose are crucial in various applications, particularly as datasets grow to substantial scales. This technical report presents an optimized parallel implementation of the Label Propagation Algorithm (LPA), a high speed community detection method, for shared memory multicore systems. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our LPA, which we term as GVE-LPA, outperforms FLPA, igraph LPA, and NetworKit LPA by 97,000×97,000$\times$97 , 000 ×, 118,000×118,000$\times$118 , 000 ×, and 40×40$\times$40 × respectively - achieving a processing rate of $1.4B$ edges/s on a $3.8B$ edge graph. In addition, GVE-LPA scales at a rate of 1.7×1.7$\times$1.7 × every doubling of threads.
@@ 1.1. Our Contributions
This report introduces GVE-LPA [^1] https://github.com/puzzlef/rak-communities-openmp, an optimized parallel implementation of LPA for shared memory multicores. On a machine with two 16-core Intel Xeon Gold 6226R processors, GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 97,000×97,000$\times$97 , 000 ×, 118,000×118,000$\times$118 , 000 ×, and 40×40$\times$40 × respectively. GVE-LPA is on average 5.4×5.4$\times$5.4 × faster than GVE-Louvain, and achieves a processing rate of $1.4B$ edges/s on a $3.8B$ edge graph. With doubling of threads, GVE-LPA exhibits an average performance scaling of 1.7×1.7$\times$1.7 ×.
@@ 5.2. Comparing Performance of GVE-LPA
We now compare the performance of GVE-LPA with FLPA (Traag and Šubelj, 2023), igraph LPA (Csardi et al., 2006), and NetworKit LPA (Staudt et al., 2016). FLPA and igraph LPA are both sequential implementations, while NetworKit LPA is parallel. For FLPA, we checkout the suitable branch with modified igraph_community_label_propagation() function, update the label propagation example in C to load the input graph from a file and measure runtime of igraph_community_label_propagation () with gettimeofday(). For igraph LPA, we follow a similar process as FLPA, but checkout code from the master branch of igraph. For NetworKit, we employ a Python script to invoke PLP (Parallel Label Propagation), and measure the total time reported with getTiming(). For each graph, we measure the NetworKit LPA runtime five times for averaging. Due to time constraints and long running times of FLPA and igraph LPA, we ran them only once per graph. Additionally, we record the modularity of communities obtained from each implementation.
Figure 4(a) shows the runtimes of FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph in the dataset, while Figure 4(b) shows the speedup of GVE-LPA with respect to each aforementioned implementation. Both FLPA and igraph LPA fail to run on kmer_A2a and kmer_V1r graphs, and hence their results are not shown. GVE-LPA exhibits an average speedup of 97,000×97,000$\times$97 , 000 ×, 118,000×118,000$\times$118 , 000 ×, and 40×40$\times$40 × over FLPA, igraph LPA, and NetworKit LPA, respectively, especially on road networks and protein k-mer graphs with low average degrees. On the sk-2005 graph, GVE-LPA identifies communities in $2.7$ seconds, achieving a processing rate of $1.4$ billion edges/s. Figure 4(c) shows the modularity of communities obtained by each implementation. On average, GVE-LPA obtains $0.6\%$ / $0.2\%$ higher modularity than FLPA and igraph LPA respectively (especially on web graphs), and $4.1\%$ lower modularity than NetworKit LPA (especially on protein k-mer graphs with large number of vertices and a low average degree).
@@ 6. Conclusion
In summary, this study concentrates on optimizing the Label Propagation Algorithm (LPA), a high-speed community detection algorithm, in the shared memory setting. We consider 6 different optimizations, which significantly improve the performance of the algorithm. Comparative assessments against competitive open-source implementations (FLPA) and packages (igraph and NetworKit) indicate that GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 97,000×97,000$\times$97 , 000 ×, 118,000×118,000$\times$118 , 000 ×, and 40×40$\times$40 ×, respectively, while identifying communities of roughly the same quality as FLPA and igraph LPA, and of $4.1\%$ lower quality (modularity) than NetworKit LPA.
@@ Acknowledgements.
I would like to thank Prof. Kishore Kothapalli, Prof. Dip Sankar Banerjee, Balavarun Pedapudi, and Souvik Karfa for their support.

### 26 Feb 2024 v4

@@ Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for this purpose are crucial in various applications, particularly as datasets grow to substantial scales. This technical report presents an optimized parallel implementation of the Label Propagation Algorithm (LPA), a high speed community detection method, for shared memory multicore systems. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our LPA, which we term as GVE-LPA, outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively - achieving a processing rate of $1.4B$ edges/s on a $3.8B$ edge graph. In addition, GVE-LPA scales at a rate of 1.7×1.7$\times$1.7 × every doubling of threads.
@@ 1.1. Our Contributions
This report introduces GVE-LPA [^1] https://github.com/puzzlef/rak-communities-openmp, an optimized parallel implementation of LPA for shared memory multicores. On a machine with two 16-core Intel Xeon Gold 6226R processors, GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively. GVE-LPA is on average 5.4×5.4$\times$5.4 × faster than GVE-Louvain, and achieves a processing rate of $1.4B$ edges/s on a $3.8B$ edge graph. With doubling of threads, GVE-LPA exhibits an average performance scaling of 1.7×1.7$\times$1.7 ×.
@@ 5.2. Comparing Performance of GVE-LPA
We now compare the performance of GVE-LPA with igraph LPA (Csardi et al., 2006), FLPA (Traag and Šubelj, 2023), and NetworKit LPA (Staudt et al., 2016). igraph LPA and FLPA are both sequential implementations, while NetworKit LPA is parallel. For FLPA, we checkout the suitable branch with modified igraph_community_label_propagation() function, update the label propagation example in C to load the input graph from a file and measure runtime of igraph_community_label_propagation () invoked with IGRAPH_LPA_FAST variant, using gettimeofday(). For igraph LPA, we follow a similar process as FLPA, but checkout code from the master branch of igraph. For NetworKit, we employ a Python script to invoke PLP (Parallel Label Propagation), and measure the total time reported with getTiming(). For each graph, we measure the NetworKit LPA runtime five times for averaging. Due to time constraints and long running times of igraph LPA, we ran it only once per graph. Additionally, we record the modularity of communities obtained from each implementation.
Figure 4(a) shows the runtimes of FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph in the dataset, while Figure 4(b) shows the speedup of GVE-LPA with respect to each aforementioned implementation. igraph LPA fails to run on protein k-mer graphs, i.e., kmer_A2a and kmer_V1r, and hence its results on these graphs are not shown. GVE-LPA exhibits an average speedup of 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × over FLPA, igraph LPA, and NetworKit LPA respectively. On the sk-2005 graph, GVE-LPA identifies communities in $2.7$ seconds, achieving a processing rate of $1.4$ billion edges/s. Figure 4(c) shows the modularity of communities obtained by each implementation. On average, GVE-LPA obtains $6.6\%$ / $0.2\%$ higher modularity than FLPA (especially on road networks) and igraph LPA respectively, and $4.1\%$ lower modularity than NetworKit LPA (especially on protein k-mer graphs).
@@ 6. Conclusion
In summary, this study focuses on optimizing the Label Propagation Algorithm (LPA), a high-speed community detection algorithm, in the shared memory setting. We consider 6 different optimizations, which significantly improve the performance of the algorithm. Comparative assessments against competitive open-source implementations (FLPA) and packages (igraph and NetworKit) indicate that GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively, while identifying communities of $6.6\%$ / $0.2\%$ higher quality than FLPA and igraph LPA respectively, and of $4.1\%$ lower quality than NetworKit LPA.
@@ Acknowledgements.
I would like to thank Prof. Kishore Kothapalli, Prof. Dip Sankar Banerjee, Balavarun Pedapudi, Souvik Karfa, and Vincent Traag for their support.